OpenStudy (anonymous):

Helpful steps please, using a sum or difference of an identity, write the following expression involving functions of x : tan(x-pi/3)?

OpenStudy (freckles):

so have you tried to apply the difference identity for tan?

OpenStudy (anonymous):

yes what I came up with was tan-sqrt(3)/1+sqrt(3) tan x, I'm unsure on my answer because it's very similar to another question and I think i'm wrong

OpenStudy (freckles):

\[\tan(x-\frac{\pi}{3})=\frac{\tan(x)-\tan(\frac{\pi}{3})}{1+\tan(x)\tan(\frac{\pi}{3})} \\ \text{ we also know } \tan(\frac{\pi}{3})= \sqrt{3} \\ \text{ so yeah you should have } \tan(x-\frac{\pi}{3})=\frac{\tan(x)-\sqrt{3}}{1+\sqrt{3} \tan(x)}\]

OpenStudy (freckles):

great job

OpenStudy (anonymous):

thanks!